OpenStudy (amonoconnor):

Determine the Anti-Derivistive of: "x^5 - 6x + 3" Any and all help is greatly appreciated! :)

OpenStudy (anonymous):

use the power rule backwards raise the power by one, divide by that number \[\int x^ndx=\frac{x^{n+1}}{n+1}\]

OpenStudy (amonoconnor):

I know, I DID the problem, and I know how to take anti's, but the answer key has something different than I got, and I'm seeking a second opinion :)

OpenStudy (anonymous):

my opinion is \[\frac{x^6}{6}-3x^2+3x\]

OpenStudy (amonoconnor):

That's close to what I got... :/ I got "(x^6)/6 - (3x^2) + 3x + C". Am I mistaken about the process of adding a "C" (representing "constant") to the answer, when dealing with Anti-Deriviatives?

OpenStudy (anonymous):

you are right, i didn't write the \(+C\)

OpenStudy (amonoconnor):

Okay. Would you mind looking at another one, falling the same scenario?

OpenStudy (amonoconnor):

It is: "1/[3(x^(1/3))]

OpenStudy (amonoconnor):

I got (x^(4/3))/4

OpenStudy (amonoconnor):

@satellite73

OpenStudy (amonoconnor):

Do you know of a different/ better answer?

OpenStudy (amonoconnor):

@satellite73 :)